OpenStudy (anonymous):

What is the sum of 7/2 + -3/5

OpenStudy (anonymous):

there is no 2.9

OpenStudy (anonymous):

29/10?

OpenStudy (anonymous):

ya thats there

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

2.9 = 29/10. One is just written as a decimal. Do you understand how to arrive at 29/10?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

Consider the denominators in your original problem. How could we arrive at a common denominator?

OpenStudy (anonymous):

what do u mean

OpenStudy (anonymous):

When we have fractions with different denominators, we cannot add or subtract straightforwardly. We first have to convert each to a common denominator. For example, if we had 1/3 + 1/4 we would convert both fractions such that each had a denominator of 12 1/3 = 4/12 1/4 = 3/12 1/3 + 1/4 = 4/12 + 3/12 = 7/12 Another way of asking the question is: what is a common multiple of 3 and 4? In this case, 12 is. So we convert each fraction so it has a denominator of 12, then add them. The same reasoning is true for your problem. What is a common multiple of 2 and 5?

OpenStudy (anonymous):

o ok thank u i get it now

OpenStudy (anonymous):

Ok! Best of luck with your studies.